# Supernovae in 2023 (review): breakthroughs by late observations

### 03 Jan 2024 v2

@@ 3.2 The breakthrough: evidence for exploding jets
In Soker (2023f) I studied the morphology of the CCSNR N63A based on new X-ray images that Karagöz et al. (2023) present. I present the image and my identification of a point-symmetric morphology in Figure 1. I identified the point-symmetric morphology to be the three pairs of ears that are marked in the figure. This point symmetry cannot be shaped by ejecta-CSM interaction or ejecta-ISM interaction (Section 1). Therefore, this point-symmetry strongly supports the JJEM. Moreover, the differences in sizes of the two ears in each pair are accounted for by the JJEM.
Figure 1: The CCSNR N63A has a point-symmetric morphology. A Chandra X-ray image of SNR N63A (logarithmic-scale $0.3-7.0{~{}\rm keV}$). The green lines and the small-yellow writings are from the original figure of Karagöz et al. (2023). The red-solid lines, the dashed-blue lines, and the red and blue writing are additions from Soker (2023f) that identify a point-symmetric morphology.
In each pair, one ear is significantly larger than the opposite ear. This is not due to an asymmetrical medium on sides A and B (marked in the figure) as Ear 1a is larger than Ear 1b, but Ears 2a and 3a are smaller than Ears 2b and 3b, respectively. I, therefore, concluded (Soker, 2023f) that the asymmetry of each pair results from asymmetrical opposite jets at the time of the jets’ launching. This asymmetry likely results from that the accretion disk that launches the jets in the JJEM has no time to fully relax. In the JJEM there are $\approx{\rm few}$ to $\approx 30$ jet-launching episodes, each lasting for $\approx 0.01-0.1{~{}\rm s}$. The viscous timescale of the accretion disk in each jet-launching episode is not much shorter than this timescale. Therefore, the disk has no time to fully relax to a thin accretion disk. Since the instabilities that must operate in the JJEM are likely to form each accretion disk with unequal sides of the temporarily equatorial plane, this asymmetry lasts for most of the jet-launching episode. The two unequal sides of the accretion disk are very likely to launch two opposite jets that have different energy, mass, and velocity. These asymmetrical two opposite jets in a pair will inflate unequal ears.
@@
SN 1987A has been one of the prime focuses of the SN community since its explosion (for 2023 papers see, e.g., Arendt et al. 2023; Cikota et al. 2023; Dohi et al. 2023; Fiorillo et al. 2023; Jones et al. 2023; Kangas et al. 2023; Larsson et al. 2023; Weishi Li et al. 2023; Sapienza et al. 2023). I am interested here in the ejecta morphology and its implication on the explosion mechanism. Ono et al. (2023) consider a bipolar explosion morphology for SN 1987A. They do not refer to jets though.
What allows me to (tentatively) claim for the JJEM for 1987A is the new observation of SN 1987A by JWST (Arendt et al. 2023) as I present in Figure 3. All marks on the original figure are from Soker (2024b), where more details can be found. In that paper, I identified a tentative point-symmetric morphology in the new JWST image of SN1987A. As said, the JJEM predicts that in some CCSNe the last jets imprint a point-symmetrical structure. I suggest such a structure in the ejecta of SN 1987A, and suggest that it was exploded by jittering jets. The asymmetry within each pair of two clumps is again attributed to the not relaxed intermittent accretion disks that launched the jets during the explosion.
@@
Let me end with the study of the northeast jet of the SNR Cassiopeia A by Koo et al. (2023b). They conclude that both the neutrino-driven explosion mechanism and the magnetocentrifugal jet-induced explosion (fixed-axis jets) have problems in accounting for the properties of the northeast jet of Cassiopeia A. This is a known result. As I commented in the past, the JJEM can best explain the properties of Cassiopeia A (e.g., Soker 2017).
I consider the point-symmetric morphologies of these three CCSNRs (and a speculative possible identification of a point-symmetric morphology in SNR CTB 1; Bear & Soker 2023), supported by earlier studies of CCSNR morphologies, to be a step toward a breakthrough in the determination of the CCSN explosion mechanism. The JJEM cannot be ignored in analyzing CCSNe.
@@ 4.1 SN Ia scenarios
The diversity of normal and peculiar SNe Ia (e.g., DerKacy et al. 2023c; Desai et al. 2023b; Dwarkadas 2023b; Fausnaugh et al. 2023; Graur et al. 2023; Harris et al. 2023; Harvey et al. 2023; Hinse et al. 2023; Holland-Ashford et al. 2023; Huang et al. 2023; Keegans et al. 2023; Larison et al. 2023; Li et al. 2023b; Lu et al. 2023; Nascimento, França, & Reis 2023; O’Brien et al. 2023; Ogawa, Maeda, & Kawabata 2023; Pearson et al. 2023b; Peterson et al. 2023; Sharon & Kushnir 2023; Singh et al. 2023; Tiwari et al. 2023; Toy et al. 2023; Xi et al. 2023; Yang et al. 2023a; Graur 2024) have motivated the consideration of all major SN Ia scenarios in 2023. I classify six scenarios as in Soker (2024a), and present them in Table 1 (for a different partition of the scenarios, in particular sub-classification of the SD scenario, see the table in the review by Liu, Röpke, & Han 2023).
@@
Group $N_{\rm exp}=1$: Lonely WD $N_{\rm exp}=2$
Outcome Mostly normal SNe Ia Mostly peculiar SNe Ia
@@
Expected in SNIPs with point-symmetric PNe. SN 1181; G1.9+0.3
Very rare: A SNIP with point-symmetric PN. SN 1181
@@
In the double degenerate (DD) scenario (e.g., Dimitriadis et al. 2023; Karthik Yadavalli et al. 2023; Moran-Fraile et al. 2023; Qi, Liu, & Wang 2023; Srivastav et al. 2023a) and in the DD-MED scenario, two WDs merge after losing angular momentum to gravitational waves. In the DD scenario, the explosion occurs within a few dynamical timescales, like in the violent merger channel, and before the non-spherical imprints of the binary system disappear. This DD scenario channel has no MED time and the explosion is non-spherical. Practically there are two stars at the explosion, i.e., $N_{\rm exp}=2$. In the DD-MED scenario channel, the explosion takes place only after the merger remnant relaxes. It is likely to be a spherical near-$M_{\rm Ch}$ explosion. At explosion, there is only one star (as wide companions play no roles), i.e., $N_{\rm exp}=1$.
The double detonation (DDet) DDet scenario, where a helium layer accreted from a companion is ignited and detonates the WD, was relatively popular in 2023 (e.g., Collins et al. 2023; Zingale et al. 2023) with attribution of this scenario to some specific peculiar SNe Ia (e.g., Liu et al. 2023b, c; Padilla Gonzalez et al. 2023a, b for four peculiar SNe Ia, respectively). Several studies attribute hyper-runaway/hypervelocity WDs to the WD helium-donor companion in the DDet scenario (e.g. El-Badry et al. 2023; Werner et al. 2023). However, new studies question whether this scenario can explain normal SNe Ia. Igoshev, Perets, & Hallakoun (2023) argue that the numbers and properties of hypervelocity WDs with velocities of $>1000{~{}\rm km}{~{}\rm s}^{-1}$ might at most account for a small fraction of SNe Ia, but might account for some peculiar SNe Ia. Braudo & Soker (2023) study the dynamics of the surviving WD and of the ejecta, and conclude that very fast runaway WDs require massive exploding CO WD. This makes such a scenario rare and supports the claim that the DDet scenario leads mostly to peculiar SNe Ia but not to normal SNe Ia.
In the single degenerate (SD) scenario (e.g., Hosseinzadeh et al. 2023b; Li, Liu, & Wang 2023, and Ruiz-Lapuente 2023 and Ruiz-Lapuente et al. 2023 for Tycho SNR) a WD accretes mass from a non-degenerate companion and explodes when reaching a near-Chandrasekhar mass. It can explode as it reaches the exploding mass, or there might be a long delay time from mass accretion to explosion (MED time). The non-detection of surviving companions in SNRs Ia, e.g., Shields et al. (2023) for SNR 0509-67.5, Tucker & Shappee (2023) for SN 2011fe, and Rodriguez, Yanza, & Dzib (2023) for Tycho’s SNR, continues to be a major challenge to the SD scenario (and also to the DDet scenario). Considering the difficulties of the SD scenario, I find it unjustified that in 2023 many papers mention only the SD and the DD scenarios, ignoring other SN Ia scenarios.
@@
Studies in 2023 have continued to explore both sub-Chandrasekhar (sub-$M_{\rm Ch}$) and near-Chandrasekhar ($M_{\rm Ch}$) explosions (e.g., Chakraborty et al. 2023; DerKacy et al. 2023a, b; Guo et al. 2023; Hoeflich et al. 2023; Leung & Nomoto 2023; Maeda et al. 2023b; Ni et al. 2023a; Yarbrough et al. 2023).
The existence of CSM around some SNe Ia continues to be puzzling (e.g. Dwarkadas 2023a; Mao-kai, Li-fan, & Xiao-feng 2023; Moriya et al. 2023b). However, it is not questionable anymore that some SNe Ia have massive CSM, as Guest et al. (2023) find for the SNR 0519–69.0 in the Large Magellanic cloud. Such is SNR G1.9+0.3 that I review in more detail in Section 4.2.3.
@@ 4.2.2 SN 1181 and its SNR Pa 30
Several papers in 2023 revealed the nature of Pa 30 as the SNR of the historical SN 1181, which was an SN Iax (Fesen, Schaefer, & Patchick, 2023; Ko et al., 2023; Lykou et al., 2023; Ritter et al., 2023). The large-scale circular morphology of Pa 30 with a hot remnant at its center is critical to the understanding of SN 1181. Fesen, Schaefer, & Patchick (2023) conduct a thorough study of Pa 30. They note that the highly spherical morphology of SNR Pa 30 is in sharp contrast to the strongly asymmetric SN Iax ejecta in some models. They attribute the filamentary structure of the round SNR to the photoionization effects of clumpy ejecta (as the modeling of some filamentary planetary nebulae). I refer here only to the large-scale spherical morphology. I, therefore, present one image from Fesen, Schaefer, & Patchick (2023) in Figure 4.

### 15 Mar 2024 v3

@@ 3.2 The breakthrough: evidence for exploding jets
In Soker (2024d) I studied the morphology of the CCSNR N63A based on new X-ray images that Karagöz et al. (2023) present. I present the image and my identification of a point-symmetric morphology in Figure 1. I identified the point-symmetric morphology to be the three pairs of ears that are marked in the figure. This point symmetry cannot be shaped by ejecta-CSM interaction or ejecta-ISM interaction (Section 1). Therefore, this point-symmetry strongly supports the JJEM. Moreover, the differences in sizes of the two ears in each pair are accounted for by the JJEM.
Figure 1: The CCSNR N63A has a point-symmetric morphology. A Chandra X-ray image of SNR N63A (logarithmic-scale $0.3-7.0{~{}\rm keV}$). The green lines and the small-yellow writings are from the original figure of Karagöz et al. (2023). The red-solid lines, the dashed-blue lines, and the red and blue writing are additions from Soker (2024d) that identify a point-symmetric morphology.
In each pair, one ear is significantly larger than the opposite ear. This is not due to an asymmetrical medium on sides A and B (marked in the figure) as Ear 1a is larger than Ear 1b, but Ears 2a and 3a are smaller than Ears 2b and 3b, respectively. I, therefore, concluded (Soker, 2024d) that the asymmetry of each pair results from asymmetrical opposite jets at the time of the jets’ launching. This asymmetry likely results from that the accretion disk that launches the jets in the JJEM has no time to fully relax. In the JJEM there are $\approx{\rm few}$ to $\approx 30$ jet-launching episodes, each lasting for $\approx 0.01-0.1{~{}\rm s}$. The viscous timescale of the accretion disk in each jet-launching episode is not much shorter than this timescale. Therefore, the disk has no time to fully relax to a thin accretion disk. Since the instabilities that must operate in the JJEM are likely to form each accretion disk with unequal sides of the temporarily equatorial plane, this asymmetry lasts for most of the jet-launching episode. The two unequal sides of the accretion disk are very likely to launch two opposite jets that have different energy, mass, and velocity. These asymmetrical two opposite jets in a pair will inflate unequal ears.
@@
SN 1987A has been one of the prime focuses of the SN community since its explosion (for 2023 papers see, e.g., Arendt et al. 2023; Cikota et al. 2023; Dohi et al. 2023; Fiorillo et al. 2023; Jones et al. 2023; Kangas et al. 2023; Larsson et al. 2023; Weishi Li et al. 2023; Bouchet et al. 2024; Chen, Chen, & Ono 2024; Fransson et al. 2024; Ravi et al. 2024; Sapienza et al. 2023). I am interested in the ejecta morphology and its implications on the explosion mechanism. Ono et al. (2023) consider a bipolar explosion morphology for SN 1987A. They do not refer to jets, though.
What allows me to (tentatively) claim for the JJEM for 1987A is the new observation of SN 1987A by JWST (Arendt et al. 2023) as I present in Figure 3. All marks on the original figure are from Soker (2024b), where more details can be found. In that paper, I identified a tentative point-symmetric morphology in the new JWST image of SN1987A. As said, the JJEM predicts that in some CCSNe, the last jets imprint a point-symmetrical structure. I suggest such a structure in the ejecta of SN 1987A and suggest that it was exploded by jittering jets. The asymmetry within each pair of two clumps is again attributed to the not relaxed intermittent accretion disks that launched the jets during the explosion.
@@
Let me end with the study of the northeast jet of the SNR Cassiopeia A by Koo et al. (2023b). They conclude that both the neutrino-driven explosion mechanism and the magnetocentrifugal jet-induced explosion (fixed-axis jets) have problems accounting for the properties of the northeast jet of Cassiopeia A. This is a known result. As I commented in the past, the JJEM can best explain the properties of Cassiopeia A (e.g., Soker 2017). New results (after this paper’s first and second versions were posted on the arXiv) by Bear & Soker (2024) show that Cassiopeia A also has a very robust point symmetry, as expected by the JJEM. An even later paper (Soker, 2024e) identifies point symmetry in more CCSNe, bringing the number to a total of eight CCSNe with point-symmetric morphology.
I consider the point-symmetric morphologies of eight CCSNRs, supported by earlier studies of CCSNR morphologies, to be a step toward a breakthrough in determining the CCSN explosion mechanism. The JJEM cannot be ignored when analyzing CCSNe.
@@ 4.1 SN Ia scenarios
The diversity of normal and peculiar SNe Ia (e.g., DerKacy et al. 2023c; Desai et al. 2023b; Dwarkadas 2023b; Fausnaugh et al. 2023; Graur et al. 2023; Harris et al. 2023; Harvey et al. 2023; Hinse et al. 2023; Holland-Ashford et al. 2023; Huang et al. 2023; Keegans et al. 2023; Larison et al. 2023; Li et al. 2023b; Lu et al. 2023; Nascimento, França, & Reis 2023; O’Brien et al. 2023; Ogawa, Maeda, & Kawabata 2023; Pearson et al. 2023b; Peterson et al. 2023; Sharon & Kushnir 2023; Singh et al. 2023; Tiwari et al. 2023; Toy et al. 2023; Xi et al. 2023; Yang et al. 2023a; Bi, Woods, & Fabbro 2024; Giuffrida et al. 2024; Graur 2024; Pakmor et al. 2024; Petrecca et al. 2024; Singh et al. 2024; Weng et al. 2024; Williams et al. 2024) have motivated the consideration of all major SN Ia scenarios in 2023. I classify six scenarios as in Soker (2024a), and present them in Table 1 (for a different partition of the scenarios, in particular sub-classification of the SD scenario, see the table in the review by Liu, Röpke, & Han 2023).
@@
[…]
@@
[…]
@@
In the double degenerate (DD) scenario (e.g., Dimitriadis et al. 2023; Karthik Yadavalli et al. 2023; Moran-Fraile et al. 2023; Qi, Liu, & Wang 2023; Srivastav et al. 2023a) and in the DD-MED scenario, two WDs merge after losing angular momentum to gravitational waves. In the DD scenario, the explosion occurs within a few dynamical timescales, like in the violent merger channel, before the binary system’s non-spherical imprints disappear. This DD scenario channel has no MED time, and the explosion is non-spherical. Practically there are two stars at the explosion, i.e., $N_{\rm exp}=2$. In the DD-MED scenario channel, the explosion occurs only after the merger remnant relaxes. It will likely be a spherical near-$M_{\rm Ch}$ explosion. At explosion, there is only one star (as wide companions play no roles), i.e., $N_{\rm exp}=1$.
The double detonation (DDet) DDet scenario, where a helium layer accreted from a companion is ignited and detonates the WD, was relatively popular in 2023 (e.g., Collins et al. 2023; Zingale et al. 2023; Boos, Townsley, & Shen 2024) with attribution of this scenario to some specific peculiar SNe Ia (e.g., Liu et al. 2023b, c; Padilla Gonzalez et al. 2023a, b for four peculiar SNe Ia, respectively). Several studies attribute hyper-runaway/hypervelocity WDs to the WD helium-donor companion in the DDet scenario (e.g. El-Badry et al. 2023; Werner et al. 2023). However, new studies question whether this scenario can explain normal SNe Ia. Igoshev, Perets, & Hallakoun (2023) argue that the numbers and properties of hypervelocity WDs with velocities of $>1000{~{}\rm km}{~{}\rm s}^{-1}$ might at most account for a small fraction of SNe Ia, but might account for some peculiar SNe Ia. Braudo & Soker (2023) study the dynamics of the surviving WD and of the ejecta, and conclude that very fast runaway WDs require massive exploding CO WD. This makes such a scenario rare and supports the claim that the DDet scenario leads mostly to peculiar SNe Ia but not to normal SNe Ia.
In the single degenerate (SD) scenario (e.g., Hosseinzadeh et al. 2023b; Li, Liu, & Wang 2023; Palicio et al. 2024, and Ruiz-Lapuente 2023 and Ruiz-Lapuente et al. 2023 for Tycho SNR) a WD accretes mass from a non-degenerate companion and explodes when reaching a near-Chandrasekhar mass. It can explode as it reaches the exploding mass, or there might be a long delay time from mass accretion to explosion (MED time). The non-detection of surviving companions in SNRs Ia, e.g., Shields et al. (2023) for SNR 0509-67.5, Tucker & Shappee (2023) for SN 2011fe, and Rodriguez, Yanza, & Dzib (2023) for Tycho’s SNR, continues to be a major challenge to the SD scenario (and also to the DDet scenario). Considering the difficulties of the SD scenario, I find it unjustified that in 2023 many papers mention only the SD and the DD scenarios, ignoring other SN Ia scenarios.
@@
Studies in 2023 have continued to explore both sub-Chandrasekhar (sub-$M_{\rm Ch}$) and near-Chandrasekhar ($M_{\rm Ch}$) explosions (e.g., Chakraborty et al. 2023; DerKacy et al. 2023a, b; Guo et al. 2023; Hoeflich et al. 2023; Leung & Nomoto 2023; Maeda et al. 2023b; Ni et al. 2023a; Yarbrough et al. 2023; Bravo, Isern, & Piersanti 2024).
The existence of CSM around some SNe Ia continues to be puzzling (e.g. Dwarkadas 2023a; Mao-kai, Li-fan, & Xiao-feng 2023; Moriya et al. 2023b; Uchida et al. 2024). However, it is not questionable anymore that some SNe Ia have massive CSM, as Guest et al. (2023) find for the SNR 0519–69.0 in the Large Magellanic cloud. Such is SNR G1.9+0.3 that I review in more detail in Section 4.2.3.
@@ 4.2.2 SN 1181 and its SNR Pa 30
Several papers in 2023 revealed the nature of Pa 30 as the SNR of the historical SN 1181, which was an SN Iax (Fesen, Schaefer, & Patchick, 2023; Ko et al., 2023; Lykou et al., 2023; Ritter et al., 2023; Ko et al., 2024). The large-scale circular morphology of Pa 30 with a hot remnant at its center is critical to the understanding of SN 1181. Fesen, Schaefer, & Patchick (2023) conduct a thorough study of Pa 30. They note that the highly spherical morphology of SNR Pa 30 is in sharp contrast to the strongly asymmetric SN Iax ejecta in some models. They attribute the filamentary structure of the round SNR to the photoionization effects of clumpy ejecta (as the modeling of some filamentary planetary nebulae). I refer here only to the large-scale spherical morphology. I, therefore, present one image from Fesen, Schaefer, & Patchick (2023) in Figure 4.